# Name That Part: 3D Part Segmentation and Naming

### v1

@@ 3.1. Architecture
Semantic Alignment: Partlet $\rightarrow$ Part Names. Each Partlet’s refined embedding $\mathbf{s}_{k}$ serves directly as its prototype $\mathbf{z}_{k}=\mathbf{s}_{k}\in\mathbb{R}^{d_{t}}$. After normalization ($\hat{\mathbf{z}}_{k}=\mathbf{z}_{k}/\|\mathbf{z}_{k}\|_{2}$), we compute cosine similarity with text embeddings $\hat{\mathbf{t}}_{a}$:
(22) $\text{sim}(k,a)=\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}$
Importantly, Partlet and text embeddings share the same semantic space ($\mathbb{R}^{d_{t}}$) without intermediate projections - this alignment is driven by the text alignment loss (Section 3.2), enabling open-vocabulary matching at inference without retraining. However, part names are often ambiguous — handle could refer to a door handle, a mug handle, or a wheelchair handle, each with distinct geometry and function. To disambiguate, we use affordance-based descriptions (e.g., “the part of a door grasped to open it” for a door handle, “the horizontal surface of a chair where a person sits” for a chair seat) generated by Gemini 2.5 Flash (comanici2025gemini). These are embedded with a sentence transformer model, such as MPNet (all-mpnet-base-v2) (song2020mpnet), for capturing long-form semantic affordances, particularly for similar part names.
3.2. Training: Partlet $\leftrightarrow$ Part Names
@@
(23) $C_{ka}=\mathcal{L}_{\text{mask}}^{(k,a)}+(1-\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a}))$
@@
Losses. Let $\mathcal{M}=\{k:\pi(k)\neq\emptyset\}$ denote matched Partlets. Our training objective combines several losses:
Text Alignment Loss. This loss is essential for open-vocabulary grounding. Without it, Partlet embeddings remain geometrically meaningful but semantically ambiguous. We apply InfoNCE contrastive loss to align matched Partlets with their text embeddings:
(24) $\mathcal{L}_{\text{text}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}-\log\frac{\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{\pi(k)}/\tau)}{\sum_{a=1}^{A}\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}/\tau)}$
@@
Mask and Partness Losses. For matched Partlets, binary cross-entropy and Dice loss supervise masks:
(25) $\mathcal{L}_{\text{mask}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}[\text{BCE}(\mathbf{m}_{k},\mathbf{m}_{\pi(k)}^{\text{gt}})$
@@
(26) $\mathcal{L}_{\text{part}}=\frac{1}{K}\sum_{k=1}^{K}\text{BCE}(\text{part}_{k},\mathbf{1}[\pi(k)\neq\emptyset])$
Auxiliary Regularizers. Coverage loss prevents over / under-segmentation by penalizing mask size disparities:
(27) $\mathcal{L}_{\text{cov}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}\left|\frac{\sum_{i}\sigma(m_{ki})-\sum_{i}m_{\pi(k)i}^{\text{gt}}}{N}\right|$
@@
(28) $\mathcal{L}_{\text{overlap}}=\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\sigma(m_{ki})-1\right)^{2}$
Global Alignment Loss. Symmetric InfoNCE aligns the global shape representation with class-level text embeddings, providing object-level semantic context:
(29) $\mathcal{L}_{\text{global}}=\frac{1}{2}\left[\mathcal{L}_{\text{CE}}(\mathbf{S},\mathbf{I})+\mathcal{L}_{\text{CE}}(\mathbf{S}^{\top},\mathbf{I})\right]$
@@
Total Loss. The complete training objective as a function of model parameters $\theta$ is:
(30) $\mathcal{L}_{\text{total}}(\theta)=$ $\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}(\theta,\pi)+\lambda_{\text{part}}\mathcal{L}_{\text{part}}(\theta,\pi)$
@@
3.3. Inference Modes and Use Cases
@@
Mode 1: Closed-Vocabulary with Confidence Calibration. We use ALIGN-Parts to scalably annotate large datasets from known categories (e.g., labeling millions of airplane meshes). This is the most practical scenario, enabling efficient 3D part segmentation and labeling with minimal oversight. We utilize this capability in building the TexParts dataset 4.7.
For training categories $\mathcal{C}$, we predict the object category via global shape-text alignment:
(31) $c^{*}=\operatorname*{arg\,max}_{c\in\mathcal{C}}\text{sim}(\mathbf{z}_{\text{global}},\mathbf{t}_{c})$
@@
(32) $C_{ka}=1-\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a})$
At inference, we use the Jonker-Volgenant algorithm (7738348) for exact optimal assignment, which is more efficient than Sinkhorn since gradients are not required.
@@
The Mahalanobis confidence (Eq. 35) requires class-conditional statistics (mean and covariance) that are estimated from the training set.
@@
(34) $\bm{\Sigma}=\text{Cov}(\{\mathbf{z}_{k}\}_{\forall k,\ell\text{ s.t. }\pi(k)=\ell})$
@@
(35) $\text{conf}_{\text{maha}}(k)=\exp(-(\mathbf{z}_{k}-\bm{\mu}_{a_{k}^{*}})^{T}\bm{\Sigma}_{a_{k}^{*}}^{-1}(\mathbf{z}_{k}-\bm{\mu}_{a_{k}^{*}}))$
@@
(36) $\text{conf}(k)=\alpha\cdot\text{conf}_{\text{soft}}(k)+(1-\alpha)\cdot\sigma(\beta\cdot(\text{conf}_{\text{maha}}(k)-0.5))$
@@
  • $\text{conf}_{\text{soft}}(k)$ is the temperature-calibrated softmax score (Eq. 38).
  • $\text{conf}_{\text{maha}}(k)$ is the Mahalanobis confidence (Eq. 35).
@@
Mode 2: Open-Vocabulary Grounding. For novel categories, users provide candidate part descriptions $\{\mathbf{t}_{a}\}_{a=1}^{A}$ (or generate via LLM from image/hint). After filtering inactive Partlets, each active Partlet matches to the best description:
(37) $a_{k}^{*}=\operatorname*{arg\,max}_{a\in\{1,\ldots,A\}}\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a})$
@@
(38) $\text{conf}_{\text{soft}}(k)=\max_{a}\frac{\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}/\tau)}{\sum_{a^{\prime}=1}^{A}\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}^{\prime}/\tau)}$
@@
Mode 3: Text-Conditioned Part Retrieval. For compatibility with Find3D (ma2024find), we retrieve a single part for query $\mathbf{t}_{q}$:
(39) $k^{*}=\operatorname*{arg\,max}_{k\in\mathcal{K}_{\text{active}}}\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{q})$
@@
(40) $\text{label}_{i}=a_{k^{*}}^{*}\quad\text{where}\quad k^{*}=\operatorname*{arg\,max}_{k\in\mathcal{K}_{\text{active}}}\sigma(m_{ki})$
@@
3.4. Implementation Details
Training Setup. ALIGN-Parts is trained on 3 NVIDIA A6000 GPUs for 2 days (batch size 16). Due to academic compute constraints, we sample 10k points per shape (vs. 100k in PartField) while maintaining strong performance. Models are normalized to $1,1]^{3}$ during training.
Architecture. The model has 34M parameters total: 5.7M for feature Fusion, 26.8M for Partlets, and 1.5M for the global classifier. Feature interactions (Partlet-to-points and Partlet-to-Partlet) use 3 transformer (vaswani2017attention) blocks with multi-head cross-attention, LayerNorm, residual connections, and feedforward layers. The BiCo fusion employs sparse 16-NN attention with 3D relative positional bias, computed via a learned MLP over Fourier-encoded (F=6 frequencies) displacement vectors, which provides geometric context while maintaining $\mathcal{O}(Nk)$ complexity. We also note that for calculating the runtime, we do not include data preprocessing time, as it varies depending on parallelization and system capabilities.
@@ 3.4. Implementation Details
Optimization. Loss weights: $\lambda_{\text{mask}}=1.0$, $\lambda_{\text{part}}=0.5$, $\lambda_{\text{text}}=1.0$, $\lambda_{\text{cov}}=0.5$, $\lambda_{\text{overlap}}=0.1$, $\lambda_{\text{global}}=1.0$. We use AdamW with an initial learning rate of 3e-4 and cosine annealing to a minimum of 5e-6.
Figure 4. Pairwise cosine similarity heatmaps between text embeddings for MPNet (left) and SigLiP (right). (Zoom in for labels)
Choice of Text Encoder We adopt MPNet (song2020mpnet) as our text encoder for part descriptions rather than CLIP / SigLIP due to its superior structure-preserving properties for sentence-level embeddings. SigLIP, optimized for short image captions (e.g., "A photo of a dog"), exhibits pathologically high cosine similarities across semantically distinct part descriptions generated by Gemini, undermining the discriminative structure necessary for partlet-based learning. In Fig.˜4 for instance, MPNet correctly assigns high similarities (>0.8) to functionally equivalent parts across classes - such as wheels (airplane, car, bicycle, wheelchair), doors (airplane, car), and handles (scooter, bicycle, wheelchair), while maintaining low similarities (<0.3) between parts with different affordances, such as tires vs. doors/windows or pedals vs. airplane components. In contrast, SigLIP assigns uniformly high similarities to both sets, collapsing the semantic space and preventing our partlets from learning meaningful text-conditioned part alignment during training.
@@
A key motivation for incorporating affordance information into part annotations is rooted in the cognitive science understanding that humans interpret and define object parts not just by geometry, but by their function, context, and description. Short or generic part names (e.g., “leg”, “handle”) are often ambiguous across different objects, lacking any semantic detail regarding the role or meaning of a part within a specific context. For example, “legs” fulfill distinct structural functions and take on different forms for chairs, tables, or sofas, a distinction that arises from their object-specific affordances. Prior work demonstrates that affordance-based cues and descriptive information guide human part recognition, reducing label ambiguity and supporting more robust reasoning and communication. Thus, by situating part annotations within functional and contextual descriptions, our approach enables higher-quality, less ambiguous labeling, consistent with cognitive models of human object understanding.
@@
Datasets. We train on 40,982 shapes from three datasets: 3DCoMPaT++ (8,627), PartNet (32,141), and Find3D (124). All use fine-grained part labels. For evaluation, we hold out 206 shapes: 126 objects (42 categories) from 3DCoMPaT++, 72 objects (24 categories) from PartNet, and 8 novel objects (8 categories) from Find3D.
4. Experiments
In this section, we discuss our methodology for unifying 3D part benchmarks with inconsistent part naming conventions for the same object classes ( Section˜4.1), introduce our baselines and metrics for the named 3D part segmentation task ( Section˜4.2), and show qualitative and quantitative comparison of ALIGN-Parts with related baselines ( Section˜4.3). We ablate different components and inference modes of ALIGN-Parts in  Section˜4.4 and  Section˜4.5. We end the section with 2 applications of our method in  Section˜4.6 and  Section˜4.7.
4.1. 3D Part Annotation Alignment & Vocabulary Compression
Our annotation alignment and vocabulary compression system employs a two-stage pipeline that combines MPNet embeddings for candidate generation and Gemini LLM verification to reject spurious matches as well as identify and merge duplicate classes and parts across a unified 3D object taxonomy. For confirmed matches, the system successfully identifies semantically equivalent entities with high MPNet similarity scores that Gemini validates as identical: for example, "laptop_computer" and "laptop" (similarity: 0.944) are merged because Gemini recognizes that "both candidate names refer to the exact same physical device… consistently define it as a portable personal computer designed for mobile use". Similarly, microwave_oven (3DCoMPaT) and microwave (PartNet) with similarity 0.902 are merged after Gemini confirmed they "describe the same kitchen appliance." Within the "microwave_oven" class, "door_glass" and "glass" (similarity: 0.865) are unified because Gemini concludes "both descriptions refer to the transparent panel integrated into the door… that allows viewing food and contains radiation. The secondary part name ’glass’ is a concise reference to the "door_glass". Similarly, "bed_footboard" and "footboard" (similarity: 0.953) are merged as Gemini states, "bed_footboard" is a more explicit naming of ’footboard’, and their descriptions are semantically identical, describing a panel at the foot of the bed opposite the headboard." For rejected pairs, the system correctly distinguishes semantically distinct parts despite high embedding similarity: "car_front_bumper" and "car_rear_bumper" (similarity: 0.879) are kept separate because Gemini determines "while both parts are bumpers with the same protective function, their specified locations (front vs. rear) make them distinct semantic parts for a 3D car object," and within the "chair" class, "back_frame_horizontal_rod" and "back_frame_vertical_rod" (similarity: 0.943) remain separate because Gemini explains "the parts are distinct based on their orientation within the back frame: one is explicitly described as a ’horizontal rod’ providing reinforcement for the backrest, while the other is a ’vertical rod’ providing structural support." The compressed vocabulary output maintains canonical names (choosing more verbose/descriptive variants), aggregates part counts across merged entities, and produces a mapping log that records every alias resolution for downstream lookup when legacy names are encountered during inference. This compressed vocabulary enables training with unified part semantics.
4.2. Comparison on Named 3D Part Segmentation
@@
Figure 5. Qualitative Results. ALIGN-Parts segments and names 3D parts robustly in a single feed-forward pass (rightmost column). Find3D (ma2024find) (first column) fails despite ground-truth part names, unable to segment the laptop in Figure 5 (bottom row). PartField (liu2025partfield) (second column) also fails: it requires ground-truth part counts for clustering, missegments bed bunks (top row), and misses refrigerator handles (second-to-last row). Our strong baseline without Partlets (third column) exhibits similar errors. In contrast, ALIGN-Parts correctly segments tiny parts, such as handles, and groups semantically similar instances (e.g., all ceiling fan blades into a single cluster).

### v2

@@ 3.1 Architecture
Semantic Alignment: Partlet $\rightarrow$ Part Names.
Each Partlet’s refined embedding $\mathbf{s}_{k}$ serves directly as its prototype $\mathbf{z}_{k}=\mathbf{s}_{k}\in\mathbb{R}^{d_{t}}$. After normalization ($\hat{\mathbf{z}}_{k}=\mathbf{z}_{k}/\|\mathbf{z}_{k}\|_{2}$), we compute cosine similarity with text embeddings $\hat{\mathbf{t}}_{a}$:
$\text{sim}(k,a)=\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}$ (7)
Importantly, Partlet and text embeddings share the same dimensional space ($\mathbb{R}^{d_{t}}$) without intermediate projections - this alignment is driven by the text alignment loss (Sec.˜3.2), enabling open-vocabulary matching at inference without retraining. However, part names are often ambiguous - handle could refer to a door handle, a mug handle, or a wheelchair handle, each with distinct geometry and function. To disambiguate, we use affordance-based descriptions (e.g., “the part of a door grasped to open it” for a door handle, “the horizontal surface of a chair where a person sits” for a chair seat) generated by Gemini 2.5 Flash [5]. These are embedded with a sentence transformer model, such as MPNet (all-mpnet-base-v2) [32], for capturing long-form semantic affordances, particularly for similar part names.
3.2 Training: Partlet $\leftrightarrow$ Part Names
@@
$C_{ka}=\mathcal{L}_{\text{mask}}^{(k,a)}+(1-\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a}))$ (8)
@@
Losses.
Let $\mathcal{M}=\{k:\pi(k)\neq\emptyset\}$ denote matched Partlets. Our training objective combines several losses:
Text Alignment Loss.
This loss is essential for open-vocabulary grounding. Without it, Partlet embeddings remain geometrically meaningful but semantically ambiguous. We apply InfoNCE contrastive loss to align matched Partlets with their text embeddings:
$\mathcal{L}_{\text{text}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}-\log\frac{\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{\pi(k)}/\tau)}{\sum_{a=1}^{A}\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}/\tau)}$ (9)
@@
Mask and Partness Losses.
For matched Partlets, binary cross-entropy and Dice loss supervise masks:
$\mathcal{L}_{\text{mask}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}[\text{BCE}(\mathbf{m}_{k},\mathbf{m}_{\pi(k)}^{\text{gt}})$ (10)
@@
$\mathcal{L}_{\text{part}}=\frac{1}{K}\sum_{k=1}^{K}\text{BCE}(\text{part}_{k},\mathbf{1}[\pi(k)\neq\emptyset])$ (11)
Auxiliary Regularizers.
Coverage loss prevents over- / under-segmentation by penalizing mask size disparities:
$\mathcal{L}_{\text{cov}}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}\left|\frac{\sum_{i}\sigma(m_{ki})-\sum_{i}m_{\pi(k)i}^{\text{gt}}}{N}\right|$ (12)
@@
$\mathcal{L}_{\text{overlap}}=\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\sigma(m_{ki})-1\right)^{2}$ (13)
Global Alignment Loss.
Symmetric InfoNCE aligns the global shape representation with class-level text embeddings, providing object-level semantic context:
$\mathcal{L}_{\text{global}}=\frac{1}{2}\left[\mathcal{L}_{\text{CE}}(\mathbf{S},\mathbf{I})+\mathcal{L}_{\text{CE}}(\mathbf{S}^{\top},\mathbf{I})\right]$ (14)
@@
Total Loss.
The complete training objective as a function of model parameters $\theta$ is:
$\mathcal{L}_{\text{total}}(\theta)=$ $\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}(\theta,\pi)+\lambda_{\text{part}}\mathcal{L}_{\text{part}}(\theta,\pi)$ (15)
@@
3.3 Inference Modes and Use Cases
@@
Mode 1: Closed-Vocabulary with Confidence Calibration. We use ALIGN-Parts to scalably annotate large datasets from known categories (e.g., labeling millions of airplane meshes). This is the most practical scenario, enabling efficient 3D part segmentation and labeling with minimal oversight. We utilize this capability in building the TexParts dataset (Sec.˜4.6). For training categories $\mathcal{C}$, we predict the object category via global shape-text alignment:
$c^{*}=\operatorname*{arg\,max}_{c\in\mathcal{C}}\text{sim}(\mathbf{z}_{\text{global}},\mathbf{t}_{c})$ (16)
@@
$C_{ka}=1-\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a})$ (17)
At inference, we use the Jonker-Volgenant algorithm [7] for exact optimal assignment, which is more efficient than Sinkhorn since gradients are not required.
@@
The Mahalanobis confidence (Eq.˜20) requires class-conditional statistics (mean and covariance) that are estimated from the training set.
@@
$\bm{\Sigma}=\text{Cov}(\{\mathbf{z}_{k}\}_{\forall k,\ell\text{ s.t. }\pi(k)=\ell})$ (19)
@@
$\text{conf}_{\text{maha}}(k)=\exp(-(\mathbf{z}_{k}-\bm{\mu}_{a_{k}^{*}})^{T}\bm{\Sigma}_{a_{k}^{*}}^{-1}(\mathbf{z}_{k}-\bm{\mu}_{a_{k}^{*}}))$ (20)
@@
$\text{conf}(k)=\alpha\cdot\text{conf}_{\text{soft}}(k)+(1-\alpha)\cdot\sigma(\beta\cdot(\text{conf}_{\text{maha}}(k)-0.5))$ (21)
@@
  • $\text{conf}_{\text{soft}}(k)$ is the temperature-calibrated softmax score (Eq.˜23).
  • $\text{conf}_{\text{maha}}(k)$ is the Mahalanobis confidence (Eq.˜20).
@@
Mode 2: Open-Vocabulary Grounding.
For novel categories, users provide candidate part descriptions $\{\mathbf{t}_{a}\}_{a=1}^{A}$ (or generate via LLM from image/hint). After filtering inactive Partlets, each active Partlet matches to the best description:
$a_{k}^{*}=\operatorname*{arg\,max}_{a\in\{1,\ldots,A\}}\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{a})$ (22)
@@
$\text{conf}_{\text{soft}}(k)=\max_{a}\frac{\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}/\tau)}{\sum_{a^{\prime}=1}^{A}\exp(\hat{\mathbf{z}}_{k}\cdot\hat{\mathbf{t}}_{a}^{\prime}/\tau)}$ (23)
@@
Mode 3: Text-Conditioned Part Retrieval.
For comparison with Find3D [26], we retrieve a single part for query $\mathbf{t}_{q}$:
$k^{*}=\operatorname*{arg\,max}_{k\in\mathcal{K}_{\text{active}}}\text{sim}(\hat{\mathbf{z}}_{k},\hat{\mathbf{t}}_{q})$ (24)
@@
$\text{label}_{i}=a_{k^{*}}^{*}\quad\text{where}\quad k^{*}=\operatorname*{arg\,max}_{k\in\mathcal{K}_{\text{active}}}\sigma(m_{ki})$ (25)
@@
3.4 Implementation Details
Training Setup.
ALIGN-Parts is trained on 3 NVIDIA A6000 GPUs for 2 days (batch size 16). Due to academic compute constraints, we sample 10k points per shape (vs. 100k in PartField) while maintaining strong performance. Models are normalized to $1,1]^{3}$ during training.
Architecture.
The model has 34M parameters total: 5.7M for feature Fusion, 26.8M for Partlets, and 1.5M for the global classifier. Feature interactions (Partlet-to-points and Partlet-to-Partlet) use 3 transformer [40] blocks with multi-head cross-attention, LayerNorm, residual connections, and feedforward layers. The BiCo fusion employs sparse 16-NN attention with 3D relative positional bias, computed via a learned MLP over Fourier-encoded (F=6 frequencies) displacement vectors, which provides geometric context while maintaining $\mathcal{O}(Nk)$ complexity. We also note that for calculating the runtime, we do not include data preprocessing time, as it varies depending on parallelization and system capabilities.
@@ 3.4 Implementation Details
Optimization.
Loss weights: $\lambda_{\text{mask}}=1.0$, $\lambda_{\text{part}}=0.5$, $\lambda_{\text{text}}=1.0$, $\lambda_{\text{cov}}=0.5$, $\lambda_{\text{overlap}}=0.1$, $\lambda_{\text{global}}=1.0$. We use AdamW with an initial learning rate of 3e-4 and cosine annealing to a minimum of 5e-6.
Figure 3: Pairwise cosine similarity heatmaps between text embeddings for MPNet (left) and SigLiP (right). (Zoom in for labels)
Choice of Text Encoder
We adopt MPNet [32] as our text encoder for part descriptions rather than CLIP / SigLIP due to its superior structure-preserving properties for sentence-level embeddings. Standard vision-language models like SigLIP often suffer from representational collapse when applied to sub-object components. Because SigLIP is optimized for holistic scene descriptions, it fails to distinguish between semantically diverse part labels (e.g., those generated by LLMs like Gemini). This results in an undifferentiated embedding space that precludes the use of discriminative partlet-based learning. In Fig.˜3 for instance, MPNet correctly assigns high similarities (>0.8) to functionally equivalent parts across classes - such as wheels (airplane, car, bicycle, wheelchair), doors (airplane, car), and handles (scooter, bicycle, wheelchair), while maintaining low similarities (<0.3) between parts with different affordances, such as tires vs. doors/windows or pedals vs. airplane components. In contrast, SigLIP assigns uniformly high similarities to both sets, collapsing the semantic space and preventing our partlets from learning meaningful text-conditioned part alignment during training.
@@
A key motivation for incorporating affordance information into part annotations is rooted in the cognitive science understanding that humans interpret and define object parts not just by geometry, but by their function, context, and description [38]. Short or generic part names (e.g., “leg”, “handle”) are often ambiguous across different objects, lacking any semantic detail regarding the role or meaning of a part within a specific context. For example, “legs” fulfill distinct structural functions and take on different forms for chairs, tables, or sofas, a distinction that arises from their object-specific affordances. Prior work demonstrates that affordance-based cues and descriptive information guide human part recognition, reducing label ambiguity and supporting more robust reasoning and communication. Thus, by situating part annotations within functional and contextual descriptions, our approach enables higher-quality, less ambiguous labeling, consistent with cognitive models of human object understanding.
Datasets.
@@
We train on 40,982 shapes from three datasets: 3DCoMPaT++ (8,627), PartNet (32,141), and Find3D (124). All use fine-grained part labels. For evaluation, we hold out 206 shapes: 126 objects (42 categories) from 3DCoMPaT++, 72 objects (24 categories) from PartNet, and 8 novel objects (8 categories) from Find3D.
@@
Figure 4: Qualitative Results. ALIGN-Parts segments and names 3D parts robustly in a single feed-forward pass (rightmost column). Find3D [26] (first column) fails despite ground-truth part names, unable to segment the laptop in Fig.˜4 (bottom row). PartField [23] (second column) also fails: it requires ground-truth part counts for clustering, missegments bed bunks (top row), and misses refrigerator handles (second-to-last row). Our strong baseline without Partlets (third column) exhibits similar errors. In contrast, ALIGN-Parts correctly segments tiny parts, such as handles, and groups semantically similar instances (e.g., all ceiling fan blades into a single cluster).
4 Experiments
In this section, we discuss our methodology for unifying 3D part benchmarks with inconsistent part naming conventions for the same object classes (Sec.˜4.1), introduce our baselines and metrics for the named 3D part segmentation task (Sec.˜4.2), and show qualitative and quantitative comparison of ALIGN-Parts with related baselines (Sec.˜4.3). We ablate different components and inference modes of ALIGN-Parts in Sec.˜4.4 and Sec.˜4.5. We end the section with 2 applications of our method in Sec.˜B.3 and Sec.˜4.6.
4.1 3D Part Annotation Alignment & Vocabulary Compression
Our 3D part annotation alignment and vocabulary compression methodology employs a two-stage pipeline that combines MPNet embeddings for candidate generation and Gemini LLM verification to reject spurious matches as well as identify and merge duplicate classes and parts across a unified 3D object taxonomy.
For confirmed matches, the system successfully identifies semantically equivalent entities with high MPNet similarity scores that Gemini validates as identical: for example, "laptop_computer" and "laptop" (similarity: 0.944) are merged because Gemini recognizes that "both candidate names refer to the exact same physical device … consistently define it as a portable personal computer designed for mobile use". Similarly, microwave_oven (3DCoMPaT) and microwave (PartNet) with similarity 0.902 are merged after Gemini confirmed they "describe the same kitchen appliance". Within the "microwave_oven" class, "door_glass" and "glass" (similarity: 0.865) are unified because Gemini concludes "both descriptions refer to the transparent panel integrated into the door … that allows viewing food and contains radiation". The secondary part name ’glass’ is a concise reference to the "door_glass". Similarly, "bed_footboard" and "footboard" (similarity: 0.953) are merged as Gemini states, " ’bed_footboard’ is a more explicit naming of ’footboard’, and their descriptions are semantically identical, describing a panel at the foot of the bed opposite the headboard."
For rejected pairs, the system correctly distinguishes semantically distinct parts despite high embedding similarity: "car_front_bumper" and "car_rear_bumper" (similarity: 0.879) are kept separate because Gemini determines "while both parts are bumpers with the same protective function, their specified locations (front vs. rear) make them distinct semantic parts for a 3D car object," and within the "chair" class, "back_frame_horizontal_rod" and "back_frame_vertical_rod" (similarity: 0.943) remain separate because Gemini explains "the parts are distinct based on their orientation within the back frame: one is explicitly described as a ‘horizontal rod’ providing reinforcement for the backrest, while the other is a ‘vertical rod’ providing structural support." The compressed vocabulary output maintains canonical names (choosing more verbose/descriptive variants), aggregates part counts across merged entities, and produces a mapping log that records every alias resolution for downstream lookup when legacy names are encountered during inference. This compressed vocabulary enables training with unified part semantics.
4.2 Comparison for Named 3D Part Segmentation
Baselines.